Find the angles of a triangle if 1 angle is 1/4 the second and the third is the sum og its two.
Answer by 304412(1)

let x be the second angle
x/4 be the ist angle
x/4 + x be the third angle
Solution:
X/4 + x + x/4 + x = 180
x+4x+x+4x/4 = 180
10x/4 = 180
10x=720
x= 72
x/4 = 72/4 = 18
x/4 + x = 18 + 72= 90
so the angles are 72, 18, 90